# Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations

### v4

@@ VI.2.2 Hellinger distance
For probabilities, their positivity $e_{i}>0$ and normalization condition $\sum_{i}e_{i}=1$ define a clove of a unitary $4-$sphere in the $\sqrt{e_{i}}-$space. Therefore, the set of nonlocal states with $|a_{1}|,|a_{2}|>|a_{3}|$ is represented in the $\sqrt{e_{i}}-$space by the intersection of the hypervolume Eq. (73) with the unitary semi-sphere centered in the origin. Since the distance in this space is the Euclidean one, the minimum in Eq. (72) in the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$ is determined by imposing the constraint
@@
$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})\equiv\displaystyle\sum_{i=1}^{4}\sqrt{e_{i}e_{i}^{loc}}+\lambda_{1}\bigg{[}(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}\bigg{]}+\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{i}^{loc}-1\right)\,.$ (75)
@@
$\begin{cases}\frac{1}{2}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0\\ \vspace{0.2cm}\par\frac{1}{2}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+2\lambda_{1}(e_{2}^{loc}-p_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{1}{2}\sqrt{\frac{e_{3}}{e_{3}^{loc}}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\par\frac{1}{2}\sqrt{\frac{e_{4}}{e_{4}^{loc}}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par\par(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{cases}$
@@
$\begin{cases}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+\sqrt{\frac{e_{3}}{e_{3}^{loc}}}=\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\,,\\ \vspace{0.2cm}\par\par\bigg{(}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}-\sqrt{\frac{e_{3}}{e_{3}^{loc}}}\bigg{)}(e_{2}^{loc}-e_{4}^{loc})=\bigg{(}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}-\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\bigg{)}(e_{1}^{loc}-e_{3}^{loc})\,,\\ \vspace{0.2cm}\par\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1\,,\\ \vspace{0.2cm}\par\par(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{cases}$
The remaining cases can be treated symmetrically; the absolute minimum is determined by the smallest value with respect to all the possible cases. In the case of Bell states, e.g. $e_{1}=1$, $e_{2}=e_{3}=e_{4}=0$, it is straightforward to verify that the minimum is reached for $e_{1}^{loc}=\frac{1+3t}{4}$, $e_{2}^{loc}=e_{3}^{loc}=e_{4}^{loc}=\frac{1-t}{4}$ , with $t=\frac{1}{\sqrt{2}}$. Therefore, as it must be, we find that the maximum value $\mathcal{M}_{He}^{max}$ of the Hellinger geometric measure of CHSH nonlocality Eq. (72) coincides with that of the Hellinger measure for Werner states, Eq. (60), with Werner parameter $w=1$.
On the other hand, we see that the above minimization procedure cannot always yield a closed formula of the Hellinger measure of CHSH nonlocality for generic Bell-diagonal states. This is implied by the type of bound on the set of local states and the form of the function one needs to minimize. Indeed, the above algebraic systems yields that in order to find the minimum one must determine the roots of a polynomial of the fifth order. Therefore, from the Abel-Ruffini-Seralian theorem it follows that it is impossible to obtain a closed analytical formula by quadratures for the roots in terms of the coefficients of the polynomial. Of course, the solutions can be determined to any desired degree of precision by standard approximation procedures such as, e.g., the Newton-Raphson method or the Laguerre iterative scheme. Alternatively, we can proceed analytically in special but relevant instances, such as the Bell states or, as we will show later on, the Bell-diagonal states with one free parameter.
VI.2.3 Bures distance
In analogy with the case of Werner states, also in the case of Bell-diagonal states the density operators commute and the Bures and Hellinger distances coincide, so that
$\mathcal{M}_{Bu}(\rho_{BD})=\mathcal{M}_{He}(\rho_{BD})\,.$ (78)
VI.2.4 Relative entropy
In the case of the relative entropy we assume that in order to determine the minimum in Eq. (14) we can continue to impose the constraint Eq. (74). In other words, we expect that also for the relative entropy of nonlocality of Bell-diagonal states the closest local state belongs to the same surface as for the Hellinger and Bures measures. Assuming the above hypothesis, the minimum in Eq. (14) is realized when
@@ VI.2.4 Relative entropy
Fixing as usual, to begin with, the case $|a_{1}|,|a_{2}|>|a_{3}|$, the Lagrange function associated to the Lagrange multipliers takes the form
$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})\equiv\displaystyle\sum_{i=1}^{4}e_{i}\log_{2}{e_{i}^{loc}}+\lambda_{1}[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}]+\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{ij}^{loc}-1\right)\,,$ (80)
@@
$\begin{cases}\frac{e_{1}}{e_{1}^{loc}\ln{2}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{p_{01}}{e_{2}^{loc}\ln{2}}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{e_{3}}{e_{3}^{loc}\ln{2}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{e_{4}}{e_{4}^{loc}\ln{2}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par(e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{cases}$
@@
$\begin{cases}\frac{e_{1}}{e_{1}^{loc}}+\frac{e_{3}}{e_{3}^{loc}}=\frac{e_{2}}{e_{2}^{loc}}+\frac{e_{4}}{e_{4}^{loc}},\\ \vspace{0.2cm}\par\bigg{(}\frac{e_{3}}{e_{3}^{loc}}-\frac{e_{1}}{e_{1}^{loc}}\bigg{)}(e_{2}^{loc}-e_{4}^{loc})=\bigg{(}\frac{e_{4}}{e_{4}^{loc}}-\frac{e_{2}}{e_{2}^{loc}}\bigg{)}(e_{1}^{loc}-e_{3}^{loc}),\\ \vspace{0.2cm}\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par(e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\par\end{cases}$
@@
In analogy with the case of the Hellinger measure, the above algebraic system is of higher order and not amenable to a solution by elementary quadratures. Therefore, in the following subsection we will consider the particular instance of one-parameter Bell-diagonal states that allows for the full analytical evaluation of the entire set of geometric and entropic measures of CHSH nonlocality.
VI.3 Bell-diagonal states with one free parameter
Let us consider Bell-diagonal states that are convex combinations of only two of the four Bell states. For instance, let us choose $e_{2}=e_{4}=0$ , $e_{1}=p$ , and $e_{3}=1-p$ , with $p$ free parameter in the interval $[0,1]$. Such Bell-diagonal states represent all possible convex combinations of the first and third Bell states of Eq. (37). Moving to symplectic space and recalling Eqs. (42), this instance corresponds to $a_{1}=a$, $a_{2}=-a$, and $a_{3}=1$, with $a=2p-1$.
For this class of states the locality condition Eq. (43) is fulfilled only for $p=\frac{1}{2}$; therefore, such states are nonlocal for $0\leq p\leq 1$, with the exception of $p=\frac{1}{2}$. The behavior of the four normalized measures of CHSH nonlocality as functions of the convex combination parameter $p$ in the half-interval $p\geq\frac{1}{2}$ is reported in Fig. 4. We see that all measures vanish at $p=\frac{1}{2}$ and reach their maximum at $p=1$, i.e. on the first Bell state. For all measures, the normalization coefficients turn out to be the same as the ones of the nonlocality measures for Werner states, in agreement with Proposition 1, Eq. (46), and Proposition 2, Eq. (53). The case $p\leq\frac{1}{2}$ is readily obtained by symmetry.
Figure 4: Normalized measures of CHSH nonlocality for the convex combination of two Bell states as functions of the convex combination parameter $p$ in the interval $p\geq\frac{1}{2}$. The same behavior holds in the symmetric case $p\leq\frac{1}{2}$, with the maximum achieved in $p=0$. Black dashed curve: Trace measure ~$\widetilde{\mathcal{M}}_{Tr}$. Black solid curve: Relative entropy measure ~$\widetilde{\mathcal{M}}_{Re}$. Grey dashed curve: Hellinger measure ~$\widetilde{\mathcal{M}}_{He}$. Grey solid curve: Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$. All plotted quantities are dimensionless.
VII Discussion and outlook
We introduced a geometric measure for Bell nonlocality in a fixed scenario, whose defining elements are the state vectors in Hilbert space and their distances from the set of local states. In the case of the CHSH inequality, we have identified the basic traits of such approach, including the set of local states and the free operations, within the framework based on Local Operations and Shared Randomness (LOSR), and we have discussed the concept of maximally nonlocal states Buscemi2012 ; Wolfe2020 ; Schmid2020b ; Rosset2020 ; Lipka2021 ; Gallego2017 .
The quantification of Bell nonlocality for quantum states in Hilbert space is naturally introduced in terms of geometric measures, in close methodological analogy with the standard theory of geometric entanglement Wei2003 ; Uyanik2010 and the more recent geometric approaches to the quantification of other types of nonlocal correlations such as, e.g., the quantum discord Geomdiscord2016 ; Geomdiscord2017 . Geometric measures define a structure and an ordering in the space of quantum states, and we have provided a detailed investigation of such structure in the case of the CHSH inequality for different classes of contractive distances (trace, Hellinger, Bures), in the case of two-qubit states, by proving and exploiting the result that for contractive distances the closest local state to a Werner state is still a Werner state, and that an analogous result holds for Bell-diagonal states. We have confronted the ordering provided by contractive geometric measures of CHSH nonlocality with the non-contractive Hilbert-Schmidt distance, and we have introduced the relative entropy of nonlocality as a further benchmark.
An important aspect of our approach is that, within a fixed Bell scenario (defined by the number of measurements, outcomes, and a Bell inequality), we perform a maximization over all possible measurement settings compatible with the given constraints. This procedure ensures that geometric measures of nonlocality depend only on the quantum state and the structure of the chosen Bell scenario, and not on the particular experimental setting. On the other hand, one could envisage an alternative strategy, in which also the measurement settings are fixed from the outset. In this case, the set of local states would be defined relative to these fixed measurements, and the resulting geometric measure would capture the nonlocality of a state specifically with respect to the selected experimental configuration, without any apriori extremization. Such a fixed-measurements approach might offer advantages in operational contexts, where the measurements are experimentally constrained or predetermined.
Generalizations and applications of the conceptual framework introduced in the present work are conceivable along different directions, such as the identification of the set of local states and the ordering of nonlocal states in finite-dimensional Hilbert spaces of higher dimension and generalized Bell inequalities, the extension to field-theoretical settings of elementary particle physics, and the use of the Hilbert space-based measures of Bell nonlocality as tools in the investigation and characterization of quantum collective phenomena and quantum phase transitions. Concerning the latter, a characterization of the ground states of quantum many-body systems in terms of Hilbert-space-based measures of Bell nonlocality could provide novel and potentially profound insights on the degree of quantumness and quantumness hierarchies between different systems of quantum matter, in analogy with and complementing the very successful use of entanglement in condensed matter theory Laflorencie2016 ; Illuminati2022 ; Illuminati2023 . Indeed, some recent preliminary studies on the structure of Bell nonlocality in quantum many-body systems hint at complex patterns of quantum nonlocality beyond entanglement Nonlocmanybody2015 ; Belldepthmanybody2019 whose precise entropic and geometric quantifications might provide important insights on the structure and hierarchy of quantum collective phenomena.
The Hilbert-space characterization and quantification that we have introduced paves the way to the study of several important questions concerning quantum nonlocality. In entanglement theory, it is known that different entanglement measures can give rise to different orderings of quantum states; for instance, the entropy-based entanglement of formation and the entanglement negativity based on the PPT criterion induce a different ordering on the set of entangled Gaussian states Adesso2005 . Similar ordering issues might arise when determining the degree of Bell nonlocality of quantum states as quantified by the different geometric or entropic measures that we have introduced in the present work. A further important question concerns the characterization and quantification of extremal nonlocality in mixed states. In entanglement theory, maximally entangled mixed states are typically defined at fixed energy, or fixed degree of local and global entropies Adesso2003 ; Adesso2004 . An investigation along similar lines could lead to the identification of the maximally nonlocal mixed states at fixed global and/or local purities.
In our opinion, a particularly challenging open problem whose eventual solution would bear far-reaching consequences revolves around the question whether the analogs of Werner states, i.e., mixed states that are local and yet entangled, exist in infinite-dimensional Hilbert spaces. In particular, it would be very interesting to determine the existence of Werner Gaussian states of two- and multimode continuous-variable systems, and to characterize their properties.
Acknowledgements.
W.R. and M.T. are supported by JST Moonshot R&D Grant No. JPMJMS2061 and JST COI-NEXT Grant No. JPMJPF2221. F.I. acknowledges funding from the Italian Ministry of University and Research, call PRIN PNRR 2022, project “Harnessing topological phases for quantum technologies”, code P202253RLY, CUP D53D23016250001, and PNRR-NQSTI project ”ECoN: End-to-end long-distance entanglement in quantum networks”, CUP J13C22000680006;

### v5

@@ B.2 Hellinger distance
For probabilities, their positivity $e_{i}>0$ and normalization condition $\sum_{i}e_{i}=1$ define a clove of a unitary 44-sphere in the $\sqrt{e_{i}}$-space. Therefore, the set of nonlocal states with $|a_{1}|,|a_{2}|>|a_{3}|$ is represented in the $\sqrt{e_{i}}$-space by the intersection of the hypervolume Eq. (77) with the unitary semi-sphere centered in the origin. Since the distance in this space is the Euclidean one, the minimum in Eq. (76) in the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$ is determined by imposing the constraint
@@
$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})$ $\equiv$ $\displaystyle\sum_{i=1}^{4}\sqrt{e_{i}e_{i}^{loc}}+\lambda_{1}\bigg[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}\bigg]$ (79)
$+$ $\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{i}^{loc}-1\right)\,.$
@@
$\left\{\begin{array}[]{l}\frac{1}{2}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \frac{1}{2}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] \frac{1}{2}\sqrt{\frac{e_{3}}{e_{3}^{loc}}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \frac{1}{2}\sqrt{\frac{e_{4}}{e_{4}^{loc}}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\[6.0pt] (e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{array}\right$. (80)
@@
$\left\{\begin{array}[]{l}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+\sqrt{\frac{e_{3}}{e_{3}^{loc}}}=\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\,,\\[6.0pt] \bigg(\sqrt{\frac{e_{1}}{e_{1}^{loc}}}-\sqrt{\frac{e_{3}}{e_{3}^{loc}}}\bigg)(e_{2}^{loc}-e_{4}^{loc})=\bigg(\sqrt{\frac{e_{2}}{e_{2}^{loc}}}-\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\bigg)(e_{1}^{loc}-e_{3}^{loc})\,,\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1\,,\\[6.0pt] (e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{array}\right$. (83)
The remaining cases can be treated symmetrically; the absolute minimum is determined by the smallest value with respect to all possible cases. In the case of Bell states, e.g. $e_{1}=1$, $e_{2}=e_{3}=e_{4}=0$, it is straightforward to verify that the minimum is reached for $e_{1}^{loc}=\frac{1+3t}{4}$, $e_{2}^{loc}=e_{3}^{loc}=e_{4}^{loc}=\frac{1-t}{4}$ , with $t=\frac{1}{\sqrt{2}}$. Therefore, as it must be, we find that the maximum value $\mathcal{M}_{\mathrm{He}}^{max}$ of the Hellinger geometric measure of CHSH nonlocality Eq. (76) coincides with that of the Hellinger measure for Werner states, Eq. (43), with the Werner parameter $w=1$.
On the other hand, we see that the above minimization procedure cannot always yield a closed formula of the Hellinger measure of CHSH nonlocality for generic Bell-diagonal states. This is implied by the type of bound on the set of local states and the form of the function that one needs to minimize. Indeed, the above algebraic systems yield that in order to find the minimum one must determine the roots of a polynomial of the fifth order. Therefore, from the Abel-Ruffini-Seralian theorem, it follows that it is impossible to obtain a closed analytical formula by quadratures for the roots in terms of the coefficients of the polynomial. Of course, the solutions can be determined to any desired degree of precision by standard approximation procedures such as, e.g., the Newton-Raphson method or the Laguerre iterative scheme. Alternatively, we can proceed analytically in special but relevant instances, such as the Bell states or the Bell-diagonal states with one free parameter.
B.3 Bures distance
In analogy with the case of Werner states, in the case of Bell-diagonal states as well, the density operators commute, and the Bures and Hellinger distances coincide, so that
$\mathcal{M}_{\mathrm{Bu}}(\rho_{\mathrm{BD}})=\mathcal{M}_{\mathrm{He}}(\rho_{\mathrm{BD}})\,.$ (84)
B.4 Relative entropy
In the case of the relative entropy, we assume that in order to determine the minimum in Eq. (20) we can continue to impose the constraint Eq. (78). In other words, we expect that for the relative entropy of nonlocality of Bell-diagonal states, the closest local state belongs to the same surface as for the Hellinger and Bures measures. Assuming the above hypothesis, the minimum in Eq. (20) is realized when
@@ B.4 Relative entropy
Fixing, as usual, to begin with the case $|a_{1}|,|a_{2}|>|a_{3}|$, the Lagrange function associated with the Lagrange multipliers takes the form
$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})$ $\equiv$ $\displaystyle\sum_{i=1}^{4}e_{i}\log_{2}{e_{i}^{loc}}+\lambda_{1}[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}]$ (86)
$+$ $\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{ij}^{loc}-1\right)\,$,
@@
$\left\{\begin{array}[]{l}\displaystyle\frac{e_{1}}{e_{1}^{loc}\ln 2}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \displaystyle\frac{p_{01}}{e_{2}^{loc}\ln 2}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] \displaystyle\frac{e_{3}}{e_{3}^{loc}\ln 2}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \displaystyle\frac{e_{4}}{e_{4}^{loc}\ln 2}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\[6.0pt] (e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{array}\right$.
@@
$\left\{\begin{array}[]{l}\displaystyle\frac{e_{1}}{e_{1}^{loc}}+\frac{e_{3}}{e_{3}^{loc}}=\frac{e_{2}}{e_{2}^{loc}}+\frac{e_{4}}{e_{4}^{loc}},\\[6.0pt] \displaystyle\bigg(\frac{e_{3}}{e_{3}^{loc}}-\frac{e_{1}}{e_{1}^{loc}}\bigg)(e_{2}^{loc}-e_{4}^{loc})=\bigg(\frac{e_{4}}{e_{4}^{loc}}-\frac{e_{2}}{e_{2}^{loc}}\bigg)(e_{1}^{loc}-e_{3}^{loc}),\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\[6.0pt] (e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{array}\right$.
@@
In analogy with the case of the Hellinger measure, the above algebraic system is of a higher order and not amenable to a solution by elementary quadratures. Therefore, in the following subsection, we will consider the particular instance of one-parameter Bell-diagonal states that allow for the full analytical evaluation of the entire set of geometric and entropic measures of CHSH nonlocality.
Appendix C Numerical Methods for the Estimation of the CHSH-Nonlocality Measures
C.1 Geometric measures
In this section, we describe the procedures used to estimate the geometric and entropic nonlocality measures for Bell-diagonal states in cases where closed analytical formulas are not available. Recall that Bell-diagonal states are convex combinations of the four Bell states and are therefore characterized by three free parameters, namely the independent probabilities $e_{1}$, $e_{2}$, and $e_{3}$, since the fourth probability is fixed by normalization as $e_{4}=1-e_{1}-e_{2}-e_{3}$. Alternatively, one can adopt the parametrization (30) in terms of the vector $\mathbf{a}=(a_{1},a_{2},a_{3})$. For the numerical analysis, the choice of parametrization can be made according to convenience, since for different distances a suitable parametrization can simplify the minimization procedure.
Symmetric families.
For Werner and isotropic states, we prove that the closest local state remains within the same family. This property allows us to derive explicit analytical expressions for the Hilbert-Schmidt, Hellinger, Bures, trace, and relative entropy measures of CHSH nonlocality (see Sections 5 and 6 of the main text). The formulas obtained for the two-qubit case are also employed as benchmarks to validate the numerical procedures described below.
Bell-diagonal states.
For general Bell-diagonal states, the minimization cannot be solved analytically, since it is not guaranteed a priori that the candidate minimum $\rho_{\mathrm{loc}}$ is a physical density operator. In particular, its positivity must be explicitly verified. The local set $\sigma$ is convex, so in principle the problem is convex once expressed in a suitable parametrization. In practice, however, the locality constraints lead to nonlinear and piecewise boundary conditions, which make direct global minimization unreliable.
To provide concrete estimation methods, we implemented a numerical strategy based on the SciPy optimization library. Specifically:
  • The parameters of Bell-diagonal states are expressed either in terms of the probabilities $e_{i}$ or the correlators $(a_{1},a_{2},a_{3})$, with explicit routines to convert between the two parameterizations.
  • Locality constraints are encoded through inequalities on the $t_{i}$ or $e_{i}$, ensuring that the candidate $\rho^{loc}\in\mathcal{L}$ corresponds to a valid density operator.
  • Since the boundary of $\sigma$ consists of disjoint surfaces, the minimization is performed separately on each surface, and the smallest value is retained. This piecewise approach improves robustness.
  • Optimization is carried out with minimize (SLSQP or trust-constr), with bounds [1,1][-1,1] on correlators and $[0,1]$ on probabilities. Multiple initial seeds (symmetric choices and random perturbations) are used to avoid local traps.
  • For entropic measures, the minimization is performed directly in the probability space, with constraints enforcing normalization and positivity.
C.2 Numerical procedure
For Bell-diagonal states, we considered two different restricted families. First, we studied convex combinations of two Bell states, which depend on a single free parameter. Second, we analyzed convex combinations of three Bell states, which involve two free parameters, with the fourth probability fixed at zero by construction. In the numerical minimization, the Hilbert-Schmidt measure is evaluated in the $(a_{1},a_{2},a_{3})$ parametrization, while for the other measures (Hellinger, Bures, trace, and relative entropy) the probability parametrization $(e_{i})$ is used. Since Werner states form a subclass of Bell-diagonal states in the two-qubit case, we use the available analytical formulas for this family as benchmarks to validate our numerical procedures. This strategy provides concrete and stable estimates of the nonlocality measures and enables the generation of contour plots over the corresponding parameter spaces.
C.3 Validation and accuracy
To ensure correctness:
  1. 1.
    We benchmark against the analytical solutions for Werner and isotropic states, reproducing the exact values within tolerance $10^{-6}$.
  2. 2.
    We check grid convergence by refining the discretization (e.g. $n=10,20,50$ points per parameter) and verifying stability of the estimates.
  3. 3.
    We verify that different initializations lead to the same minimum, confirming consistency of the optimization.
All Python codes used for these computations are available on GitHub: https://github.com/2zanfardino/geometric_nonlocality. Typical execution times for two-parameter contour plots (grid size $50\times 50$) are a few minutes on a standard workstation.